Fermi National Laboratory


Accelerator Update for March 2, 2005 - March 4, 2005

Wednesday March 2
The day shift began with Linac down due to work on the H- Source. This keeps all beam off.

A Linac expert reported at 9:22 AM that the H- Source was back online.

Operations resumed sending beam to the Antiproton Source (Pbar) at 9:38 AM.

Operators prepared the Meson enclosures M01 through M05 for access at 9:57 AM.

Operations put the Tevatron (TeV) through a dry squeeze at 1:10 PM.

Operations put the TeV into shot setup at 1:41 PM.

At 2:18 PM, a vacuum leak in the TeV sector F11, P2 line, put the shot setup on hold.

At 4:46 PM, The TeV suffered a sector D1 quench just before the store was to be established.

Operations made the Collider Detector at Fermilab (CDF) and D-Zero (D0) ready for access at 5:19 PM.

Vacuum technicians accessed F sector at 5:39 PM to find and repair the vacuum problem.

The vacuum techs reported at 8:51 PM that they had found the F sector vacuum leak. By 9:16 PM, they had successfully patched the beam pipe.

CDF came out of access at 9:38 PM.

Operations established beam to MiniBooNE at 9:58 PM.

Operations established beam to the Main Injector (MI) and to Pbar by 10:06 PM.

Operations put the TeV through a dry squeeze at 11:21 PM.

At 11:39 PM, Operations reported that a MI RF station (MIRF7) started tripping off and would not stay on when reset. Operators bypassed station.

Thursday March 3
The midnight shift began with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE taking beam.

Operations loaded a 12x0 study-store into the TeV at 12:42 AM.

At 4:30 AM, Operations reported that Pbar had been suffering from an emittance blow up. Operations reloaded a data file to resolve the problem.

At 6:10 AM, the TeV quenched during studies due to an abort kicker prefire.

The Neutron Therapy Facility (NTF) requested beam at 9:03 AM for patient treatment, which would occur around 10:30 AM.

At 9:33 AM, Operations reported that Drift Tube #1 in Linac had tripped off for the first of four times in about six minutes. A Linac expert investigated and discovered a Linac quadrupole power supply at full scale. They adjusted the output and stacking resumed at 10:03 AM

Operations put the TeV into shot setup at 11:47 AM.

At 2:10 PM, Operations reported that about 1/3 of the antiprotons fell out of MI during transfers.

Operations established store 4021 at 2:49 PM with an initial luminosity of 74.04E30.

Pbar resumed stacking at 3:01 PM.

Operations reported at 3:26 PM that NuMI was taking beam.

At 4:13 PM, a Linac expert reported that a RF station (LRF4) was sparking and needed about ten minutes of work.

At 5:49 PM, a Linac RF station (KRF5) tripped off for the first of three times in an hour. It reset okay each time.

At 6:06 PM, LRF3 tripped off for the first of three times in six minutes. It reset okay each time.

At 6:55 PM, KRF5 tripped off. It reset okay.

Friday March 4
The midnight shift began with Operations monitoring store 4021, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 1:25 AM, Operations reported that MIRF7 had tripped off for the first of seven times in six hours. It reset okay each time.

NuMI studies finished for the night at 3:17 AM.

Operations put the Recycler into shot setup at 3:17 AM to stash antiprotons.

Recycler transfers finished and Pbar resumed stacking at 5:15 AM.

Operations established beam to the Meson experiment MCenter at 6:39 AM.

The Plan for Friday and the Weekend
The overall plan for today and the weekend is to stack, stash, and store. The specific plan for today is to monitor the store until the NuMI commissioning event occurs, around 3:15 PM, and then to allow TeV experts begin an end-of-store study around 3:30 PM. Operations will put the TeV into shot setup around 4:30 PM.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/4/2005   email Fermilab